write the standard equation of the parabola with the following characteristics
focus (-2,1)
directrix x=3
Answer by DSMLMD(16) 
 Focus point at (-2,1) and directrix x = 3
directrix
x = a - p
3 = a - p .... (1)
focus point
F((p+a),b)
F(-2,1)
b = 1
p + a = -2 .... (2)
eliminate (1) and (2) we get
p = -2.5
a = 0.5
because the directrix located at x-coordinate, the parabola equation is:
(y-b)^2 = 4p(x-a)
(y-1)^2 = 4(-2.5)(x-0.5)
(y-1)^2 = -10(x-0.5)
so, the parabola equation is (y-1)^2 = -10(x-0.5)
|
|
|
| |